OpenStudy (anonymous):

simplfy 2 radical 112

OpenStudy (anonymous):

8 root 7

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

\[112=16\times 7\] \[\sqrt{16\times 7}=\sqrt{16}\times \sqrt{7}=4\sqrt{7}\]

OpenStudy (anonymous):

\[2\sqrt{112}=8\sqrt{7}\]